Circular No. 6640

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1997 H1 (McNAUGHT-HUGHES) J. V. Scotti, Lunar and Planetary Laboratory; and A. Nakamura,

Kuma Kogen Astronomical Observatory, have independently recovered comet P/1991 S1 (= 1991y = 1991 IX) as shown below. The indicated correction to the prediction by the undersigned on MPC 25183 (ephemeris on MPC 27943) is ΔT = +0.02 day.

1997 UT α_{2000} δ_{2000} m_{2} Observer
Apr. 16.44773 15^{h}22^{m}59.^{s}38 −10^{o}31′03′′.0 20.6 Scotti
16.47023 15 22 58.57 −10 31 00.0 20.6  _{”}
16.48596 15 22 58.04 −10 30 58.0 20.1  _{”}
29.65512 15 14 06.69 − 9 58 52.3 20.3 Nakamura
29.66510 15 14 06.18 − 9 58 50.8  _{”}
29.67682 15 14 05.69 − 9 58 48.6  _{”}

J. V. Scotti and J. Larsen (Kitt Peak). 0.9-m Spacewatch telescope. A. Nakamura (Kuma Kogen). 0.60-m f/6.0 Ritchey-Chrétien. Comet es-

sentially stellar in appearance. Communicated by S. Nakano.

Improved orbital elements from 33 observations 1991–1997:

Epoch = 1998 Mar. 8.0 TT

T = 1998 Feb. 23.7633 TT  ω = 224.^{o}3657
e = 0.404125  Ω = 89.9731 }2000.0
q = 2.116237 AU  i = 7.3033

a = 3.551479 AU n^{o} = 0.1472619 P = 6.693 years

SUPERNOVA 1997by IN ANONYMOUS GALAXY M. Turatto and S. Benetti, European Southern Observatory, report:

“We obtained a spectrogram (range 320–1000 nm, resolution 1.2 nm) on Apr. 30.18 UT with the 1.5-m telescope at La Silla. The spectrum of the new supernova closely resembles that of the prototypal SN 1994D a couple of days before maximum, thus indicating that SN 1997by is a normal type Ia. Strong absorption lines superimposed on a blue continuum are identified as due to Si II, S II, Fe III and Ca II. The redshift of the parent galaxy Abell 1736 (measured at the SN position using Na I, Mg I and the G band) is z = 0.045. This redshift implies a photospheric velocity (deduced from the Si II line) of 10 760 km/s for SN 1997by, also similar to the value for SN 1994D.”

1997 April 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden